OpenStudy (anonymous):

Write the first five terms of each sequence. a) an=4n-1 b) an=3(n+1) c) an=3a n-1-1,a1=2 d) an= 2a n-1-n,a1=6

OpenStudy (jmark):

Given a) an=4n-1 a_1=4-1 a_1=3 b) an=3(n+1) a_1=3(1+1) a_1=3*2 a_1=6 c) an=3an-1-1,a1=2 a_2=3(2)-1 a_2=6-1 a_2=5 d) an= 2a n-1-n,a1=6 a_2=2(6)-2 a_2=12-2 a_2=10